1/X + 1/Y = -1/Z Y=2 and z=3
1/X + 1/2 = -1/3
- 1/2 -1/2
1/X = -5/6
X = -5/6
Please help is this correct???
Thanks


Answer by stanbon(75887)

1/X + 1/Y = -1/Z Y=2 and z=3
1/X + 1/2 = -1/3
- 1/2 -1/2
1/X = -5/6
Invert both sides to solve for X:

X = -6/5
================
Cheers,
Stan H.